# Updating GUT-Scale Pole Higgs Inflation After ACT

### v1

@@ Abstract
Abstract: We consider models of chaotic inflation driven by the real parts of a conjugate pair of Higgs superfields involved in the spontaneous breaking of a grand unification symmetry at a scale assuming its value within MSSM. We combine a superpotential, which is uniquely determined by applying a continuous $R$ symmetry, with two fractional shift-symmetric Kähler potentials introducing two free parameters $(p,N)$. The inflationary observables provide an excellent match to the recent ACT data for $1.355\leq p\leq 6.7$ and $6\cdot 10^{-5}\leq N\leq 0.7$. The attainment of inflation allows for subplanckian inflaton values and possibly detectable primordial gravitational waves with $(p,N)$ values of order unity. A solution to the $\mu$ problem of MSSM and baryogenesis via non-thermal leptogenesis can be also accommodated extending the superpotential of the model with suitable terms.
@@ I Introduction
The announcement of Data Release 6 (DR6) from the Atacama Cosmology Telescope (ACT) act ; actin fuelled a plethora of works act0 ; act2 ; act4 ; gup ; oxf ; indi ; kina ; actlee ; rhc ; rhb ; rha ; act5 ; nmact ; maity ; reh8 ; actattr ; act1 ; actj ; yin ; actpal ; act3 ; act6 ; actellis ; acttamv ; ketov ; r2a ; r2b ; r2drees ; r2mans ; r2li ; heavy ; fhi1 ; fhi2 ; fhi3 ; fhi4 ; fhi5 which revise many well-motivated inflationary models – for a review see Ref. review . This is because the combination of the aforementioned data with the cosmic microwave background (CMB measurements by Planck plin and BICEP/Keck (BK) bcp , together with the Dark Energy Spectroscopic Instrument (DESI) Baryon Acoustic Oscillation (BAO) results desi suggests a value of the (scalar) spectral index $n_{\rm s}$ substantially larger than the one indicated from the Planck data plin . Namely, the so-called P-ACT-LB-BK18 data entails actin
@@
Below, in Sec. II, we describe the building blocks ($W$ and $K$’s) of our proposal, we outline the derivation of the inflationary potential in Sec. III and test our model against observations in Sec. IV. Finally, we expose a post-inflationary completion in Sec. V and summarize our conclusions in Sec. VI. Unless otherwise stated, we use units where the reduced Planck scale $m_{\rm P}=2.43\cdot 10^{18}~{\mbox{\rm GeV}}$ is equal to unity.
@@ II Model Set-up
The starting point of our approach is the superpotential known from the models of F-term hybrid inflation fhi1 ; fhi2 ; fhi3 ; fhi4 ; fhi5
@@
which is uniquely determined, at renormalizable level, by a gauge symmetry $\mathbb{G}$ and a convenient continuous $R$ symmetry. Here, $\lambda$ and $M$ are two constants which can both be taken positive by field redefinitions; $S$ is a left-handed superfield, singlet under $\mathbb{G}$; $\bar{\Phi}$ and $\Phi$ is a pair of left-handed superfields belonging to non-trivial conjugate representations of $\mathbb{G}$, and reducing its rank by their vacuum expectation values (v.e.vs). Just for definiteness we restrict ourselves to $\mathbb{G}_{B-L}=\mathbb{G}_{\rm SM}\times U(1)_{B-L}$ which consists the simplest GUT beyond the MSSM – here $\mathbb{G}_{\rm SM}$ is the gauge group of MSSM and $B$ and $L$ denote the baryon and lepton number. With the specific choice of $\mathbb{G}$ the $U(1)_{B-L}$ and $R$ charges of the various superfields are
$(B-L)(S,\bar{\Phi},\Phi)=(0,-2,2)~~~\mbox{and}~~~R(S,\bar{\Phi},\Phi)=(2,0,0).$ (3)
At the SUSY limit $W$ leads to a BLB-L phase transition since the vacuum of the model lies at the direction
$\langle{S}\rangle\simeq 0~~\mbox{and}~~\langle{\Phi}\rangle=\langle{\bar{\Phi}}\rangle\simeq M/\sqrt{2}~~\mbox{for}~~M\ll 1,$ (4)
i.e., $U(1)_{B-L}$ is spontaneously broken via $\langle{\Phi}\rangle$ and $\langle{\bar{\Phi}}\rangle$.
@@ III Inflationary Potential
The appropriateness of the proposed $W$ and $\widetilde{K}_{\ell\rm st}$ in Eqs. (2) and (5) can be verified if we identify the inflationary potential and the canonically normalized inflaton with those introduced in Ref. actpole within SUGRA.
To embark on it, we focus on the lagrangian density for the complex scalar fields $z^{\alpha}=S,\Phi,\bar{\Phi}$ – denoted by the same superfield symbol – which can be written as linde1
@@
where $z^{\alpha}$ are complex scalar fields – the symbol $z^{\alpha}$ as subscript denotes derivation with respect to (w.r.t) $z^{\alpha}$ – and $D_{\mu}$ is the gauge covariant derivative. Also $V$ is the SUGRA potential which can be found in terms of $W$ and $K$’s in Eqs. (2) and (5) via the formula
@@
To specify the canonically normalized inflaton, we note that, for both $\widetilde{K}_{\ell\rm st}$’s in Eq. (5), $K_{{\alpha}{\bar{\beta}}}$ along the configuration in Eq. (14) takes the form
@@ 44 Weyl $\widehat{\psi}_{\pm}$ $m^{2}_{\psi\pm}$ $6f_{\rm T}^{p+2}H_{\rm HI}^{2}/pN\phi^{2}f_{p}$}
We can verify that the inflationary direction in Eq. (14) is stable w.r.t the fluctuations of the fields $\chi^{\alpha}=\theta_{+},\theta_{\Phi}$ and $S$. To this end, we construct the mass-squared spectrum of those scalars. Taking the limit $\phi\sim 1$ we find the expressions of the masses squared $m^{2}_{\chi^{\alpha}}$ arranged in Table 1, which approach rather well the exact expressions taken into account in our numerical computation. These expressions assist us to appreciate the role of $N_{\rm st}>0$ in retaining positive $m^{2}_{S}$. Moreover, $m^{2}_{\chi^{\alpha}}\gg H_{\rm HI}^{2}=V_{\rm I0}/3$ for $\phi_{\rm f}\leq\phi\leq\phi_{\star}$ – where $\phi_{\star}$ and $\phi_{\rm f}$ are the values of $\phi$ when $k_{\star}=0.05/{\rm Mpc}$ crosses the horizon of ppHI and at the end of ppHI correspondingly. In Table 1 we display also the masses, $M_{BL}$, of the gauge boson $A_{BL}$ and the masses of the corresponding fermions with the eigenstate $\widehat{\psi}_{\pm}$ defined as $\widehat{\psi}_{\pm}=(\widehat{\psi}_{\Phi+}\pm{\psi}_{S})/\sqrt{2}$. Since $\mathbb{G}_{B-L}$ is broken during ppHI, no cosmic strings are formed after its termination.
@@ IV Inflationary Requirements
${N_{\star}}\simeq 61.5+\ln{V_{\rm HI}(\phi_{\star})^{\frac{1}{2}}\over V_{\rm HI}(\phi_{\rm f})^{\frac{1}{4}}}~~\mbox{and}~~A_{\rm s}^{\frac{1}{2}}\simeq 4.618\cdot 10^{-5},$ (20)
where we consider an equation-of-state parameter $w_{\rm rh}\simeq 1/3$ correspoding to quartic potential which is expected to approximate rather well $V_{\rm HI}$ for $\phi\ll 1$. The constraints above allows us to extract $\lambda$ and $\phi_{\star}$ for any given $p$, NN and $M$. However, $M$ is irrelevant for the inflationary observables, provided that it is confined to values much lower than $m_{\rm P}$. Motivated by the unification of the gauge coupling constants with the particle content of MSSM, we determine $M$ identifying the unification scale $M_{\rm GUT}\simeq 2/2.433\cdot 10^{-2}$ with the value of $M_{BL}$ – see Table 1 – at the SUSY vacuum in Eq. (4). Given that $\langle{f_{\rm T}}\rangle\simeq 1$ for $M\ll 1$, we obtain, for both $\widetilde{K}_{\ell\rm st}$ in Eq. (5)
@@
which are indepedent of the exponent $n=4$ of $\phi$ in Eq. (15) whereas $r$ for $n=4$ is written as
@@
The expressions above for $n_{\rm s}$ and $a_{\rm s}$ are independent from $N$ in the lowest order expansion in powers of $1/{N_{\star}}$. This feuture, however, is not totally kept as we show numerically in Fig. 1, where we compare our outputs against the P-ACT-LB-BK18 data actin in the $n_{\rm s}-r$ plane. We draw solid, dashed, dot-dashed and dotted lines for $p=1,2,5$ and $10$ respectively and show the variation of $N$ along each line. For the two latter $p$ values $N_{\rm min}$’s from Eq. (22) are also displayed. We clearly see that $n_{\rm s}$ increases with $p$. However, no upper bound on $p$ can be inferred from the upper bound on $n_{\rm s}$ in Eq. (1) and therefore we take $p=10$ as the maximal $p$ value. It is evident that the whole observationally favored range in the $n_{\rm s}-r$ plane is covered varying $p$ and $N$. As a bonus, $r$ values for $(p,N)$’s of order unity lie within the reach of planned experiments det aiming to discover primordial gravitational waves.
Fig. 1: Allowed curves in the $n_{\rm s}-r$ plane for various NN’s indicated on the curves. The conventions for the selected $p$ values are shown in the legends. The marginalized joint $68\%$ $95\%$] regions from P-ACT-LB-BK18 data are depicted by the dark [light] shaded contours.
Our plot in Fig. 1 reveals that $n_{\rm s}$ slightly decreases as $N$ increases and so we can obtain a lower bound on the pNp-N plane. The relevant allowed region is shown in Fig. 2. It is bounded by (i) the dashed line which originates the bound on $r$ in Eq. (1), (ii) the dot-dashed line which comes from the lower bound on $n_{\rm s}$ in Eq. (1) and (iii) the dotted line along which Eq. (22) is saturated. Fixing $n_{\rm s}$ to its central observational value in Eq. (1), we obtain the solid line along which we obtain the following allowed ranges
@@
with ${N_{\star}}\simeq(54.7-56.8)$, $a_{\rm s}\simeq-4.7\cdot 10^{-4}$ and $r\gtrsim 2.8\cdot 10^{-4}$. The obtained $|a_{\rm s}|$’s remain negligibly small being, thereby, consistent with its $95\%$ observationally allowed margin actin .
@@
Fig. 2: Allowed (shaded) region as determined by Eq. (20) in the pNp-N plane. The conventions adopted for the various lines are also shown.
@@ V Post-Inflationary Completion
$\Delta K$ $=$ $\mbox{$\sum_{a=1}^{5}$}N_{\rm st}\ln\left(1+|X_{\alpha}|^{2}/N_{\rm st}\right)$ (25b)
with  $X_{\alpha}=N^{c}_{i},L_{j},{H_{u}}$ and $H_{d}$}. Here we adopt the notation and the BLB-L and $R$ charges of the various superfields as displayed in Table 1 of Ref. nmBL . For simplicity we use the same prefactors of $\ln$ in all terms of $\Delta K$.
@@
$\langle{{H_{u}}}\rangle_{\rm I}=\langle{{H_{d}}}\rangle_{\rm I}=\langle{\widetilde{N}^{c}_{i}}\rangle_{\rm I}=0,$ (26)
where $\widetilde{N}^{c}_{i}$ are the scalar superpartners of the right-handed neutrinos $N^{c}_{i}$. The stability of this path can be checked, parameterizing the complex fields above as we do for $S$ in Eq. (13) – note that we use small letters for the real and imaginary components. The relevant masses squared are listed in Table 2 where we see that $M^{2}_{iN^{c}}>0$ and $m^{2}_{h+}>0$ for any $\phi_{\rm f}<\phi<1$ – here $m_{h\pm}$ are the eigenvalues associated with the Higgs eigenstates
@@
We can verify numerically that the requirement above is fulfilled for $N_{\rm st}=1$ and $\lambda_{\mu}\lesssim 2\cdot 10^{-5}$. Despite this low upper bound, $\lambda_{\mu}$ can become compatible with $\mu/m_{3/2}\sim 1$ as we show in Sec. A, where the generation of the $\mu$ term of MSSM is analyzed. On the other hand, the implementation of nTL is explored in Sec. B and our final results, including the post-inflationary constraints, are given in Sec. C. Note in passing that due to the extra contributions of Table 2 $\Delta V_{\rm HI}$ increases and so $\Lambda$ should be readjusted. E.g., for $p=2$ we find $\Lambda\simeq(4-5)\cdot 10^{15}~{\mbox{\rm GeV}}$ – hereafter we restore units, i.e., we take $m_{\rm P}=2.433\cdot 10^{18}~{\mbox{\rm GeV}}$.
Table 2: Mass-squared spectrum of the non-inflaton sector along the path in Eqs. (14) and (26) for $K=\widetilde{K}_{\ell\rm st}+\Delta K$.
@@
states $\widetilde{K}_{\rm 1st}+\Delta K K=K~$$\widetilde{K}_{\rm 2st}+\Delta K$
@@
Scalars $\tilde{\nu}^{c}_{i},\bar{\tilde{\nu}}^{c}_{i}$ $m_{i\tilde{\nu}^{c}}^{2}$ $3H_{\rm HI}^{2}\left(1+1/N_{\rm st}+8\lambda^{2}_{iN^{c}}\phi^{2}/\lambda^{2}\phi^{2}\right)$

### v2

@@ Abstract
Abstract: We consider models of chaotic inflation driven by the real parts of a conjugate pair of Higgs superfields involved in the spontaneous breaking of a grand unification symmetry at a scale assuming its value within MSSM. We combine a superpotential, which is uniquely determined by applying a continuous $R$ symmetry, with two fractional shift-symmetric Kähler potentials introducing two free parameters $(p,N)$. The inflationary observables provide an excellent match to the recent ACT data for $1.355\leq p\leq 6.7$ and $6\cdot 10^{-5}\leq N\leq 0.7$. The attainment of inflation allows for subplanckian inflaton values and possibly detectable primordial gravitational waves with $(p,N)$ values of order unity. A solution to the $\mu$ problem of MSSM and baryogenesis via non-thermal leptogenesis can be also accommodated by embedding the model into a BLB-L extension of MSSM.
@@ I Introduction
The announcement of Data Release 6 (DR6) from the Atacama Cosmology Telescope (ACT) act ; actin fuelled a plethora of works act0 ; act2 ; act4 ; gup ; oxf ; indi ; kina ; actlee ; rhc ; rhb ; rha ; act5 ; nmact ; maity ; reh8 ; actattr ; act1 ; actj ; yin ; actpal ; act3 ; act6 ; act7 ; act8 ; actellis ; acttamv ; ketov ; r2a ; r2b ; r2drees ; r2mans ; r2li ; heavy1 ; heavy ; fhi1 ; fhi2 ; fhi3 ; fhi4 ; fhi5 ; smth which revise many well-motivated inflationary models – for a review see Ref. review . This is because the combination of the aforementioned data with the cosmic microwave background (CMB measurements by Planck plin and BICEP/Keck (BK) bcp , together with the Dark Energy Spectroscopic Instrument (DESI) Baryon Acoustic Oscillation (BAO) results desi suggests a value of the (scalar) spectral index $n_{\rm s}$ substantially larger than the one indicated from the Planck data plin . Namely, taking into account also the tensor-to-scalar ratio $r$, the so-called P-ACT-LB-BK18 data entails actin
@@
Below, in Sec. II, we describe the building blocks ($W$ and $K$’s) of our proposal, we outline the derivation of the inflationary potential in Sec. III and test our model against observations in Sec. IV. Finally, we expose a post-inflationary completion in Sec. V and summarize our conclusions in Sec. VI. In Appendix A we present our approach to the reheating stage in our model. Unless otherwise stated, we use units where the reduced Planck scale $m_{\rm P}=2.43\cdot 10^{18}~{\mbox{\rm GeV}}$ is equal to unity.
@@ II Model Set-up
Our starting point is the superpotential known from the models of F-term hybrid inflation fhi1 ; fhi2 ; fhi3 ; fhi4 ; fhi5
@@
which is uniquely determined, at renormalizable level, by a gauge symmetry $\mathbb{G}$ and a convenient continuous $R$ symmetry. Here, $\lambda$ and $M$ are two constants which can both be taken positive by field redefinitions; $S$ is a left-handed superfield, singlet under $\mathbb{G}$; $\bar{\Phi}$ and $\Phi$ is a pair of left-handed superfields belonging to non-trivial conjugate representations of $\mathbb{G}$, and reducing its rank by their vacuum expectation values (v.e.vs). Just for definiteness we restrict ourselves to $\mathbb{G}_{B-L}=\mathbb{G}_{\rm SM}\times U(1)_{B-L}$ which consists the simplest GUT beyond the MSSM – here $\mathbb{G}_{\rm SM}$ is the gauge group of MSSM and $B$ and $L$ denote the baryon and lepton number. The BLB-L and $R$ charges of $S,\bar{\Phi}$ and $\Phi$ can be assigned as shown in Table 1 of Ref. nmBL . At the SUSY limit $W$ leads to a BLB-L phase transition since, at the SUSY vacuum
$\langle{S}\rangle\simeq 0~~\mbox{and}~~\langle{\Phi}\rangle=\langle{\bar{\Phi}}\rangle=M/\sqrt{2}~~\mbox{for}~~M\ll 1,$ (3)
$U(1)_{B-L}$ is spontaneously broken via $\langle{\Phi}\rangle$ and $\langle{\bar{\Phi}}\rangle$.
@@ III Inflationary Potential
The appropriateness of the proposed $W$ and $\widetilde{K}_{\ell\rm st}$ in Eqs. (2) and (4) can be verified if we identify the inflationary potential and the canonically normalized inflaton with those introduced in Ref. actpole for Tp-Model inflation.
To embark on it, we focus on the SUGRA lagrangian density for the complex scalar fields $z^{\alpha}=S,\Phi,\bar{\Phi}$ – denoted by the same superfield symbol – which can be written as linde1
@@
where $z^{\alpha}$ are complex scalar fields – the symbol $z^{\alpha}$ as subscript denotes derivation with respect to (w.r.t) $z^{\alpha}$ – and $D_{\mu}$ is the gauge covariant derivative. Also $V$ is the SUGRA potential which can be found in terms of $W$ and $\widetilde{K}_{\ell\rm st}$ in Eqs. (2) and (4) via the formula
@@
To specify the canonically normalized inflaton, we note that, for both $\widetilde{K}_{\ell\rm st}$ in Eq. (4), $K_{{\alpha}{\bar{\beta}}}$ along the configuration in Eq. (13) takes the form
@@ 44 Weyl $\widehat{\psi}_{\pm}$ $m^{2}_{\psi\pm}$ $6f_{\rm T}^{p+2}H_{\rm HI}^{2}/pN\phi^{2}f_{p}$}
We can verify that the inflationary direction in Eq. (13) is stable w.r.t the fluctuations of the fields $\chi^{\alpha}=\theta_{+},\theta_{\Phi}$ and $S$. To this end, we construct the mass-squared spectrum of those scalars. Taking the limit $M\ll 1$ we find the expressions of the masses squared $m^{2}_{\chi^{\alpha}}$ arranged in Table 1, which approach rather well the exact expressions taken into account in our numerical computation. These expressions assist us to appreciate the role of $N_{\rm st}>0$ in retaining positive $m^{2}_{S}$. Moreover, $m^{2}_{\chi^{\alpha}}\gg H_{\rm HI}^{2}=V_{\rm I0}/3$ for $\phi_{\rm f}\leq\phi\leq\phi_{\star}$ – where $\phi_{\star}$ and $\phi_{\rm f}$ are the values of $\phi$ when $k_{\star}=0.05/{\rm Mpc}$ crosses the horizon of ppHI and at the end of ppHI correspondingly. In Table 1 we display also the masses, $M_{BL}$, of the gauge boson $A_{BL}$ and the masses of the corresponding fermions with the eigenstate $\widehat{\psi}_{\pm}$ defined as $\widehat{\psi}_{\pm}=(\widehat{\psi}_{\Phi+}\pm{\psi}_{S})/\sqrt{2}$. Since $\mathbb{G}_{B-L}$ is broken during ppHI, no cosmic strings are formed after its termination.
@@ IV Inflationary Requirements
${N_{\star}}\simeq 61.3+\frac{1-3w_{\rm rh}}{12(1+w_{\rm rh})}\ln\frac{\pi^{2}g_{\rm rh*}T_{\rm rh}^{4}}{30V_{\rm HI}(\phi_{\rm f})}+$
$\frac{1}{4}\ln{V_{\rm HI}(\phi_{\star})^{2}\over g_{\rm rh*}^{1/3}V_{\rm HI}(\phi_{\rm f})}~~\mbox{and}~~A_{\rm s}^{\frac{1}{2}}\simeq 4.618\cdot 10^{-5},$ (19)
where we take into account that ppHI is followed in turn by an oscillatory phase with mean equation-of-state parameter $w_{\rm rh}$, radiation and matter domination. We estimate $w_{\rm rh}$ as detailed in Appendix A and we find $w_{\rm rh}\simeq(0.12-0.3)$ in the major part of the parameter space – note that $w_{\rm rh}$ decreases with $N$.
The constraints above allows us to extract $\lambda$ and $\phi_{\star}$ for any given $p$, NN and $M$. However, $M$ is irrelevant for the inflationary observables, provided that it is confined to values much lower than $m_{\rm P}$. Motivated by the unification of the gauge coupling constants with the particle content of MSSM, we determine $M$ identifying the unification scale $M_{\rm GUT}\simeq 2/2.433\cdot 10^{-2}$ with the value of $M_{BL}$ – see Table 1 – at the SUSY vacuum in Eq. (3). Given that $\langle{f_{\rm T}}\rangle\simeq 1$ for $M\ll 1$, we obtain, for both $\widetilde{K}_{\ell\rm st}$ in Eq. (4)
@@
which are independent of the exponent $n=4$ of $\phi$ in Eq. (14) whereas $r$ for the specific $n$ value is written as
@@
The expressions above for $n_{\rm s}$ and $a_{\rm s}$ are independent from $N$ in the lowest order expansion in powers of $1/{N_{\star}}$. This feature, however, is not totally kept as we show numerically in Fig. 1, where we compare our outputs against the P-ACT-LB-BK18 data actin in the $n_{\rm s}-r$ plane. We draw solid, dashed, dot-dashed and dotted lines for $p=1,2,5$ and $10$ respectively and show the variation of $N$ along each line. For the two latter $p$’s the corresponding $N_{\rm min}$’s from Eq. (21) are also displayed. We clearly see that $n_{\rm s}$ increases with $p$ and $r$ with $N$ and so the whole shaded region favored by P-ACT-LB-BK18 is covered varying $p$ and $N$. As a bonus, the resulting $r$’s for $(p,N)$’s of order unity lie within the reach of planned experiments det aiming to discover primordial gravitational waves.
Fig. 1: Allowed curves in the $n_{\rm s}-r$ plane for various $p$’s shown in the legend and varying NN as shown along the curves. The marginalized joint $68\%$ $95\%$] regions from P-ACT-LB-BK18 data are depicted by the dark [light] shaded contours.
The variation of the observables in Fig. 1 reveals that our free parameters can be constrained in the pNp-N plane. The relevant allowed region is shown in Fig. 2. It is bounded by (i) the dashed line which originates the bound on $r$ in Eq. (1), (ii) the dot-dashed line which comes from the lower bound on $n_{\rm s}$ in Eq. (1) and (iii) the dotted line along which Eq. (21) is saturated – we take by hand a maximal $p=10$ since no upper bound on $p$ can be inferred from the corresponding bound on $n_{\rm s}$ in Eq. (1). Fixing $n_{\rm s}$ to its central value in Eq. (1), we obtain the solid line along which we obtain
@@
with ${N_{\star}}\simeq(54.7-56.8)$, $a_{\rm s}\simeq-4.7\cdot 10^{-4}$ and $r\gtrsim 2.8\cdot 10^{-4}$. The obtained $|a_{\rm s}|$’s remain negligibly small being, thereby, consistent with its $95\%$ c.l. allowed margin in Ref. actin .
@@
Fig. 2: Allowed (shaded) region as determined by Eqs. (1), (IV) and (21) in the pNp-N plane. Along the solid line we fix $n_{\rm s}=0.974$. The conventions adopted for the various lines are also shown.
@@ V Post-Inflationary Completion
$\Delta K$ $=$ $\mbox{$\sum_{a=1}^{4}$}N_{\rm st}\ln\left(1+|X_{\alpha}|^{2}/N_{\rm st}\right)$ (24b)
with  $X_{\alpha}=\tilde{N}^{c}_{i},\tilde{L}_{j},{H_{u}}$ and $H_{d}$} where $\widetilde{N}^{c}_{i}$ and $\tilde{L}_{j}$ are the scalar superpartners of the right-handed neutrinos $N^{c}_{i}$ and left-handed leptons $L_{j}$ respectively. Here we adopt the notation and the BLB-L and $R$ charges of the various superfields as displayed in Table 1 of Ref. nmBL . For simplicity we use the same prefactors of $\ln$ in all terms of $\Delta K$.
@@
$\langle{{H_{u}}}\rangle_{\rm I}=\langle{{H_{d}}}\rangle_{\rm I}=\langle{\widetilde{N}^{c}_{i}}\rangle_{\rm I}=\langle{\tilde{L}_{j}}\rangle_{\rm I}=0.$ (25)
The stability of this path can be checked, parameterizing the complex fields above as we do for $S$ in Eq. (12) – note that we use small letters for the real and imaginary components. The relevant masses squared are listed in Table 2 where we see that $m_{i\tilde{\nu}^{c}}^{2}>0$, $m_{i\tilde{l}}^{2}>0$ and $m^{2}_{h+}>0$ for any $\phi_{\rm f}<\phi<1$ – here $m_{h\pm}$ are the eigenvalues associated with the Higgs eigenstates
@@
We can verify numerically that the requirement above is fulfilled for $N_{\rm st}=1$ and $\lambda_{\mu}\lesssim 2\cdot 10^{-5}$. Despite this low upper bound, $\lambda_{\mu}$ can become compatible with $\mu/m_{3/2}\sim 1$ as we show in Sec. A, where the generation of the $\mu$ term of MSSM is analyzed. On the other hand, the implementation of nTL is explored in Sec. B and our final results, including the post-inflationary constraints, are given in Sec. C. Note in passing that due to the extra contributions of Table 2 $\Delta V_{\rm HI}$ increases and so $\Lambda$ should be readjusted. E.g., for $p=2$ we find $\Lambda\simeq(4-5)\cdot 10^{15}~{\mbox{\rm GeV}}$ – hereafter we restore units, i.e., we take $m_{\rm P}=2.43\cdot 10^{18}~{\mbox{\rm GeV}}$.
Table 2: Mass-squared spectrum of the non-inflaton sector along the path in Eqs. (13) and (25) for $K=\widetilde{K}_{\ell\rm st}+\Delta K$ with $\ell=1,2$.
@@
states
@@
Scalars $\tilde{\nu}^{c}_{i},\bar{\tilde{\nu}}^{c}_{i}$ $m_{i\tilde{\nu}^{c}}^{2}$ $3H_{\rm HI}^{2}\left(1+1/N_{\rm st}+8\lambda^{2}_{iN^{c}}/\lambda^{2}\phi^{2}\right)$
$\tilde{l}_{i},\bar{\tilde{l}}_{i}$ $m_{i\tilde{l}}^{2}$ $3H_{\rm HI}^{2}(1+1/N_{\rm st})$